# Languages of Boundedly-Ambiguous Vector Addition Systems with States

### v1

@@ 6 Deciding regularity
We devote the rest of this section to the proof of Theorem 6.1. Let us fix a $k$-ambiguous VASS $V$. We show, how to decide regularity of a language of a $k$-ambiguous VASS. Using Theorem 6.2 we get a downward-VASS $\hat{V}$ recognising the complement of $L(V)$. Observe, that $L(V)$ is regular if and only if $L(V)$ and $L(\hat{V})$ are regular separable. Now we will use the following claim about downward-VASSs.
@@
We obtain reachability VASS $V^{\prime}$ recognising the same language as $V$ by taking VASS $V$ and for each $q(u)\in D$ adding state $q_{q(u)}$, $\varepsilon$ transition from with no effect from $q$ to $q_{q(u)}$. Moreover, for each $i\in[1,d]$ if $u_{i}\in\mathbb{N}$ we add $\varepsilon$ transition from $q_{q(u)}$ to $q_{q(u)}$ incrementing by one the $i$th counter and otherwise if $u_{i}=\omega$ we add an $\varepsilon$-transition from $q_{q(u)}$ to $q_{q(u)}$ decrementing by one $i$th counter. Let the set of added states be equal to $Q^{\prime}$. We set the initial configurations of $V^{\prime}$ to be the initial conditions of $V$. For vector $u\in(\mathbb{N}\cup\omega)^{d}$ let us denote by $\hat{u}\in\mathbb{N}^{d}$ vector such that for all $i\in[1,d]$ we have $u_{i}=\hat{u}_{i}$ if $u_{i}\in\mathbb{N}$ and $\hat{u}_{i}=0$ otherwise. We set the accepting configurations of $V^{\prime}$ to the following set of configurations $F^{\prime}=\{q_{q(u)}(\hat{u})\mid q_{q(u)}\in Q^{\prime}\}$. Now we have to prove, that $L(V)=L(V^{\prime})$. For $L(V)\subseteq L(V^{\prime})$ observe, that for each $w\in L(V)$ there exists configuration $q(v)\in F$ such that $w$ is read by an accepting run $\rho$ ending in $q(v)$. As $F$ is a downward-closed set there is $u$ such that $v\preceq u$ such that $q_{q(v)}(\hat{v})\in F^{\prime}$. Moreover, observe that $q_{q(v)}(\hat{v})$ is reachable from $q(u)$ in $V$ using $\varepsilon$-transitions. Hence $w\in L(V^{\prime})$.
@@ 7 Future research
Other problems for boundedly-ambiguous VASSs. We have shown in Section 6 that the regularity problem is decidable for baVASSs (boundedly-ambiguous VASSs), in contrast to general VASSs. It is also known that the language equivalence problem is decidable for baVASSs [DBLP:journals/lmcs/CzerwinskiH25], while being undecidable for general VASSs [DBLP:journals/tcs/ArakiK76, DBLP:conf/lics/HofmanMT13]. It is natural to ask whether other classical problems are decidable for baVASSs, for example deciding whether there exists an equivalent deterministic or unambiguous VASS. These problems are undecidable for general VASS due to Theorem 1.1, but can possibly be decidable for baVASSs. One can also ask whether given $k$-ambiguous VASS has an equivalent $(k-1)$-ambiguous VASS, our techniques from Section 5 used for showing undecidability does not seem to work in that case. Other further research for baVASSs would be to ask about complexity of the mentioned problems, for example to understand complexity of the regularity problem for baVASSs, since we already know it is decidable.
@@ 8 Acknowledgements
We would like to thank David Purser for sharing with us a conjecture about language not recognized by an unambiguous VASS. We thank also Sławomir Lasota for suggesting us to reduce from 0-finite-reach instead of reducing from the halting problem for 2CM in the proof of Theorem 1.1. Additionally, we would like to thank Piotr Hofman for inspiring discussions and for reviewing the Master’s Thesis [Orlikowski2024], which served as a foundation for the results presented in this paper.

### v2

@@ 6 Deciding regularity
We devote the rest of this section to the proof of Theorem 6.1. Let us fix a $k$-ambiguous VASS $V$. We show how to decide regularity of a language of a $k$-ambiguous VASS. Using Theorem 6.2 we get a downward-VASS $\hat{V}$ recognising the complement of $L(V)$. Observe, that $L(V)$ is regular if and only if $L(V)$ and L(V^)L($\hat${V}) are regular separable. Now we will use the following claim about downward-VASSs.
@@
[…]
@@ 7 Future research
Other problems for boundedly-ambiguous VASSs. We have shown in Section 6 that the regularity problem is decidable for baVASSs (boundedly-ambiguous VASSs), in contrast to general VASSs. It is also known that the language equivalence problem is decidable for baVASSs [DBLP:journals/lmcs/CzerwinskiH25], while being undecidable for general VASSs [DBLP:journals/tcs/ArakiK76, DBLP:conf/lics/HofmanMT13]. It is natural to ask whether other classical problems are decidable for baVASSs, for example deciding whether there exists an equivalent deterministic or unambiguous VASS. These problems are undecidable for general VASS due to Theorem 1.1, but can possibly be decidable for baVASSs. One can also ask whether given $k$-ambiguous VASS has an equivalent (k1)(k-1)-ambiguous VASS, our techniques from Section 5 used for showing undecidability does not seem to work in that case. Other further research for baVASSs would be to ask about complexity of the mentioned problems, for example to understand complexity of the regularity problem for baVASSs, since we already know it is decidable.
@@ 8 Acknowledgements
We would like to thank David Purser for sharing with us a conjecture about language not recognised by an unambiguous VASS. We thank also Sławomir Lasota for suggesting us to reduce from 0-finite-reach instead of reducing from the halting problem for 2CM in the proof of Theorem 1.1. Additionally, we would like to thank Piotr Hofman for inspiring discussions and for reviewing the Master’s Thesis [Orlikowski2024], which served as the foundation for the results presented in this paper.